Hanging Greens, Inc., buys hanging plants for $4.50 each and resells them for $7.95 each. The firm sells 18,000 plants per year. The firm orders 600 plants at a time and has a fixed cost per order of $39. The carrying cost per unit is $.80. To avoid newer plants mixing with older plants, the inventory is totally sold out before it is restocked. The total annual carrying cost is ____ and the total annual restocking cost is
Answer #1

Total annual carrying cost = [(600 + 0)/2] × $.80 = $240

Total restocking cost = (18,000/600) × $39 = $1,170

The total annual carrying cost is $240 and the total annual restocking cost is $1,170.